OpenStudy (sherribelle):

I'm doing inverse functions. If f(x) and f^-1(x) are inverse functions of each other and f(x) = 2x + 5, what is f^-1(8)? A. -1 B. 3/2 C. 41/8 D. 23

OpenStudy (sherribelle):

I've done an equation like this but since it comes out as a fraction I'm not sure on what to do.

OpenStudy (anonymous):

what is the inverse?

OpenStudy (freckles):

So y=f(x) and f^{-1}(y)=x So basically you find when 2x+5=8

OpenStudy (freckles):

Are if you want you can find the inverse first and then plug in directly into the inverse.

OpenStudy (freckles):

But solving 2x+5=8 will suffice.

OpenStudy (akashdeepdeb):

f(x) = 2x + 5 OR y = 2x + 5 Now, to find the inverse, we have to represent x in terms of y, and then switch the variables x and y. [This is a method to find the inverse of any function] STEP 1 : Represent x in terms of y. y = 2x + 5 y - 5 = 2x x = (y-5)/2 STEP 2 : Switch the variables. y = (x-5)/2 \(f^{-1} (x) = \frac{(x-5)}{2}\) Now find f^{-1} (8). :)

OpenStudy (anonymous):

use @AkashdeepDeb method. That is appropriate.

OpenStudy (sherribelle):

So we're substituting x for 8? Sooooo 8 -5 = 3 divided by 2 equals 3/2

OpenStudy (anonymous):

yes, you got it

OpenStudy (freckles):

2x+5=8 also gives you that.

OpenStudy (anonymous):

it does but theoretically, you didn't state the inverse of f(x).

OpenStudy (sherribelle):

You're both right, but if you're trying to teach someone that doesn't quite understand it you would use the other method. Even though your method is simple.

OpenStudy (anonymous):

In a test, you would be given points to show the approach of the inverse and the substitution.

OpenStudy (freckles):

Theoretically you do not need to find the actually inverse function to find the inverse at a certain number. :)